Need help with Algebra word problem.
A science teacher has a 2 pint solution of 3% sulfuric acid. She needs to get a solution of 4% sulfuric acid. She only has a solution of 10% sulfuric acid to add to the 3% sulfuric acid. How much 10% sulfuric acid would she have to add to the 3% sulfuric acid to get the needed 4% solution.

Answer by Paul(988)

10x+3(2)=4(x+2)
10x+6=4x+8
6x=2
x=1/3

Hence, about 1/3 solution should be added.
Paul.